OpenStudy (anonymous):

given f(x)=3x-1 and g(x)=2x-3, for which value of x does g(x)=f(2)?

OpenStudy (usukidoll):

so we need to figure out what makes g(x) = f(2) so we are given that for f(x) x = 2 so plug in x = 2 into f(x) = 3x-1 and see what number we have.. then we can find an x and plug it into g(x) to make g(x) = f(2) true

OpenStudy (anonymous):

so f(x) would = 7

OpenStudy (usukidoll):

wrong arithmetic. If f(x) = 3x+1 x = 2 f(2) = 3(2)+1=6+1=7 but we are given f(x) = 3x-1 so plug x = 2 in the correct function .

OpenStudy (usukidoll):

so f(2) = 3(2) -1 ? we are subtracting

OpenStudy (usukidoll):

@tildax what is 6 - 1?

OpenStudy (anonymous):

sorry just got back lol and 5 haha

OpenStudy (usukidoll):

it's alright lol so f(2) = 5 so now we need an x for g(x) to produce 5 g(x) = f(2) g(x) =5

OpenStudy (usukidoll):

g(x)=2x-3 what x value do we need to produce 5?

OpenStudy (anonymous):

4

OpenStudy (usukidoll):

alright! so when x = 4 g(4) = 2(4)-3 =8-3=5 so g(4) = f(2) 5=5 :)

OpenStudy (anonymous):

yay tyty

OpenStudy (anonymous):

how do i medal you or what ever for thanks haha

OpenStudy (usukidoll):

click on best response button

OpenStudy (anonymous):

i did :)

OpenStudy (usukidoll):

yay I got it :)